# AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels

### v1

@@ 3.4. Self-Learning Generator
After generating the query, we employ the generator to produce a hypothetical document that encapsulates the correct information based on the true target document. We intentionally avoid using the true target document as the output label because the primary role of the generator is to craft a hypothetical document that aids the retriever in locating the true target document. Expecting the generator to replicate the exact target document itself would be overly demanding and unrealistic.
(6) $\{d_{1}^{\prime},...d_{N}^{\prime}\}=\mathcal{M}_{g}(q,d,Prompt_{3})$
@@
Considering that the hypothetical documents generated by the generator may not always be ideal for retrieving the true target document. Ideally, the optimal hypothetical document is one that enhances the ranking position of the target document during retrieval. To achieve this, we employ the retriever $\mathcal{M}_{r}$ to rank and select the optimal hypothetical document. Specifically, the generator creates $K$ hypothetical documents for a given query, and each hypothetical document is utilized by the retriever to search the corpus for the most relevant documents. We then calculate the ranking position $r_{i}$ of the true target document for $d^{\prime}_{i}$:
(7) $r_{i}=rank(d,\mathcal{M}_{r}(d^{\prime}_{i},D)),i=1,...,K$
(8) $d^{*}=d^{\prime}_{\mathrm{arg\ min^{K}_{i=1}}\ r_{i}}$
The higher the ranking, the more effective the hypothetical document. We select the generated document with the highest rank (i.e., the lowest $r_{i}$). At this stage, we have constructed a question-answer pair in the form of $(q,d^{*})$, where the query serves as the question, and the generated document acts as the answer. We then employ a supervised fine-tuning approach to train the generator using this dataset, $D_{llm}=\{(q,d^{*})|q\in Q\}$. The standard supervised fine-tuning (SFT) loss is calculated as follows:
(9) $\mathcal{L}_{\mathrm{sft}}=-\sum\nolimits_{q\in Q}\sum\nolimits_{t}\mathrm{log}\ \mathcal{M}_{g}(q^{\prime}_{t}|q^{\prime}_{\textless t},q)$
For the training data of the self-learning generator, there is no need to rely on supervision signals from labeled medical data. Instead, we utilize only unlabeled corpora, leveraging the generator and retriever within the SL-HyDE framework to create the supervision signals. More importantly, the self-learning generator trained in this manner can produce hypothetical documents optimized for retrieval ranking preferences, thereby enhancing the retriever’s performance in subsequent retrieval processes.
3.5. Self-Learning Retriever
The training of retrievers typically involves two stages: unsupervised pre-training and supervised fine-tuning. Unsupervised pre-training aims to harness large volumes of unlabeled text to learn domain knowledge, while supervised fine-tuning employs high-quality labeled datasets, often alongside hard negative mining, to further teach the model to differentiate between similar texts. For example, GTE (Li et al., 2023) utilizes $\sim$800M text pairs during the unsupervised pre-training stage and $\sim$3M pairs for fine-tuning. In this work, our goal is not to train a medical retriever with vast amounts of data but rather to adapt a general-purpose retriever to the medical domain through fine-tuning with a small amount of data.
[…]
In addition to the unlabeled medical corpus, we now have a corresponding query. This $(q,d)$ combination already forms the query-document labeled data required for retriever training. However, considering that HyDE retrieves documents by simultaneously encoding both the query and the hypothetical document to locate the target document, we propose using a $(q,d^{\prime},d)$ triplet as the labeled data for retriever training. This method narrows the gap between the training and inference stages for the retriever, ensuring the two stages operate under the same data format.
@@ 3.5. Self-Learning Retriever
To achieve this, we utilize the fine-tuned generator $\mathcal{M}_{g}$ from the previous stage to generate hypothetical documents for all queries, constructing a labeled fine-tuning dataset $D_{emb}=\{(q,d^{\prime},d>)\}$. Following previous research (Xiao et al., 2024; Li et al., [n. d.]), we enhance the complexity of the training data through hard negative mining using the retriever, which creates challenging negative examples that closely resemble the queries. The hard negative dataset $D^{-}$ is mined from the original corpus $D$, employing the ANN-style sampling strategy in (Xiong et al., 2020):
(10) $D^{-}=ANN(\mathcal{M}_{r}(q,d^{\prime}),\mathcal{M}_{r}(d))$
@@
In addition to the negatives mined from the corpus, we also incorporate in-batch negatives. We then apply contrastive learning loss for supervised fine-tuning of the retriever, with the objective function formulated as follows:
(11) $\mathcal{L}_{\mathrm{emb}}=\mathrm{min.}\sum_{(q,d)}-\mathrm{log}\ \frac{e^{s(q,d)/\tau}}{e^{s(q,d)/\tau}+\sum\nolimits_{B\cup D^{-}}e^{s(q,d^{-})/\tau}}$
@@
where $\tau$ is the temperature coefficient, and $B$ represents the negative samples within the batch. The similarity score $s(q,d)$ incorporates the generated document, as described in Equation 4.
Thus far, we have developed a retriever equipped with medical domain knowledge that is coherently adapted to the characteristics of retrieval queries, incorporating hypothetical documents within the HyDE framework.
3.6. SL-HyDE vs. HyDE
Our approach, SL-HyDE, builds upon HyDE with several enhancements while retaining some similarities. First, both SL-HyDE and HyDE follow the same inference process for information retrieval tasks. Each uses a large model to generate a hypothetical document based on the query, which the retriever then employs to locate the most relevant document. Second, neither SL-HyDE nor HyDE requires labeled data, which allows for rapid deployment. HyDE is especially advantageous in real-world scenarios where efficient retrieval can be executed simply by selecting a generator and a retriever. However, for tasks needing domain-specific knowledge, such as medical information retrieval, deploying HyDE directly may not yield optimal results. One potential strategy is to fine-tune the generator and retriever separately using labeled medical data before deploying the HyDE framework. The primary challenge here in acquiring labeled data, and fine-tuning the models separately often leads to suboptimal performance.
SL-HyDE improves upon this by integrating a self-learning mechanism, transforming HyDE into a trainable end-to-end framework. This mechanism enables both the generator and the retriever to better adapt to the medical domain without relying on labeled medical data. Supervision signals for the generator’s training are derived from the retriever, and vice versa, facilitating mutual enhancement through this self-learning process. This holistic approach results in improved performance in retrieval tasks. Overall, SL-HyDE offers an efficient and convenient solution for enhancing HyDE’s performance in the medical domain, particularly when dealing with unlabeled corpora.
4. CMIRB BENCHMARK
The CMIRB benchmark is a specialized multi-task dataset designed specifically for medical information retrieval. Our collection and construction methodology is guided by the following principles: (i) Domain Specificity: The benchmark focuses exclusively on medical tasks, providing a targeted evaluation setting. (ii) Task Diversity: It includes a variety of query types and document formats to reflect real-world medical scenarios. (iii) Task Difficulty: The tasks are crafted to challenge existing models, thereby pushing the boundaries of medical retrieval research.
Figure 2. An overview of the diverse tasks and datasets in CMIRB benchmark.
4.1. Task Definition
Figure 2 provides an overview of the tasks and datasets available in CMIRB. The benchmark comprises the following five task types:
Medical Knowledge Retrieval :    Retrieve relevant medical knowledge snippets from textbooks or encyclopedias based on a given medical entity query.
Medical Consultation Retrieval :     Extract relevant doctor’s responses to online medical consultation questions posed by patients, simulating the retrieval of expert medical advice.
Medical News Retrieval :     Focus on retrieving news articles that address queries related to COVID-19.
Medical Post Retrieval :    Retrieve the content of a forum post corresponding to its title.
Medical Literature Retrieval :     Retrieve abstracts of cited references based on the title of a medical paper or find a similar paper based on the text of a given medical paper.
4.2. Data Construction
The data corpus for CMIRB is partially sourced from existing datasets and supplemented by high-quality data collected from relevant medical websites.
For the medical knowledge retrieval task, we develop three datasets: the medical exam retrieval dataset MedExam, the medical encyclopedia knowledge retrieval dataset DuBaike, and the medical disease knowledge retrieval dataset DXYDisease. For MedExam, the queries and corpus are derived from the MEDQA(MCMLE) test dataset and MedTextbook(MCMLE) (Jin et al., 2021). We concatenate the multiple-choice question and options as the query. For DuBaike, the queries are derived from the DuReader dataset (He et al., 2017), while the documents consist of medical knowledge pages from Baidu Baike [^3] https://baike.baidu.com/. These queries are generated by users while searching on Baidu Search or Baidu Zhidao. For DXYDisease, we collect question-document pairs from the online disease knowledge website DingXiangYuan [^4] https://dxy.com/. The questions consist of pre-constructed inquiries about various aspects of diseases, such as definitions and symptoms.
For the medical consultation retrieval task, we collect three online patient-doctor consultation datasets: MedicalRetrieval, CmedqaRetrieval, and DXYConsult. The MedicalRetrieval dataset is sourced from (Long et al., 2022), and the CmedqaRetrieval dataset comes from (Qiu et al., 2022). Additionally, we compile several doctor-patient dialogues from the DingXiangYuan, to create DXYConsult. Compared to the former two datasets, the query in DXYConsult includes more detailed patient information, such as symptoms, medication usage, and pending diagnosis questions.
For the medical news retrieval task, we utilize the CovidRetrieval dataset from (Qiu et al., 2022). This dataset features manually annotated high-quality questions related to COVID-19 as queries, with corresponding news articles as the documents.
In the medical post retrieval task, we curate IIYiPost by crawling posts from the IIYi forum [^5] https://bbs.iiyi.com/, a platform where medical professionals discuss various medical topics. This dataset includes both post titles and the full post content.
@@ 4.2. Data Construction
For the medical literature retrieval task, we develop two datasets based on the CSL data (Li et al., 2022): the citation retrieval dataset CSLCite and similar literature retrieval dataset CSLRel. To construct CSLCite, we utilize journal titles from CSL as queries and extract their cited references from WanFangMedical [^6] https://med.wanfangdata.com.cn/ as positive documents. Similarly, we select the most similar paper recommended by the website as the positive documents for the CLSRel dataset.
To ensure dataset quality, we design a filtering process, as outlined in Algorithm 1. We use ChatGPT to exclude non-medical data and remove low-quality query-document pairs. For the MedExam and DuBaike datasets, beyond the initial filtering, we implement a query-document matching process to find the most similar positive document for each query, detailed in Appendix A.1. After processing, we compile a total of 10 datasets. The statistical details of these datasets are presented in Table 1. Statistics reveal that query lengths vary significantly, ranging from a single title to a full article. Similarly, document lengths span from brief physician responses to extensive medical knowledge texts. This diversity ensures that our benchmark is both comprehensive and practically meaningful.
Table 1. Statistics of datasets in CMIRB.
@@
Table 2. Performance of various Retrieval models on CMIRB benchmark. All scores denote NDCG@10. The best score among 10 base retrievers on a given dataset is marked in bold.
@@
5. EXPERIMENTS
5.1. Experimental Setup
Implementation Details    For the training medical corpus, we sample 10,000 documents from the Huatuo26_encyclopedia dataset. In our training framework, we utilize Qwen2-7B-Instruct (Yang et al., 2024) as the generator and BGE-Large-zh-v1.5 (Xiao et al., 2024) as the retriever. Unless otherwise stated, all subsequent experiments employ the Qwen+BGE setup. Model training and evaluation are conducted on up to 5 NVIDIA A100 GPUs with 40GB of memory each. During fine-tuning of the LLM, the AdamW optimizer (Loshchilov, 2017) is employed alongside a cosine-type scheduler. Training is executed for 1 epoch with a learning rate of 1e-5 and a batch size of 2. The warmup steps for the LLM are set to 200, and the lora rank is configured at 8. For retriever training, we use the AdamW optimizer with an initial learning rate of 1e-5, applying a linear decay schedule. The batch size per GPU is set at 4, and the maximum input sequence length is limited to 512. The model is trained for 1 epoch. We set the temperature to 0.02 and mined 7 hard negatives for training.
Evaluation Settings   For simplicity, we employ the LLM to generate a single hypothetical document for each query. The retrieval model embeds all queries, hypothetical documents, and corpus documents, with similarity scores calculated using cosine similarity. Documents in the corpus are ranked for each query based on these scores, and nDCG@10 is used as the primary evaluation metric to assess retrieval effectiveness.
Baseline Models   To comprehensively evaluate CMIRB, we compare our model against several popular retrieval models. These include lexical retriever BM25 (Robertson et al., 2009); dense retrieval models such as Text2Vec-Large-Chinese [^7] https://github.com/chmlyrical/text2vec-large-chinese, PEG (Wu et al., 2023), BGE-Large-zh-v1.5 (Xiao et al., 2024), GTE-Large-zh (Li et al., [n. d.]), and Piccolo-Large-zh [^8] https://github.com/timczm/piccolo-large-zh; multilingual retrievers like mContriever (masmarco) (Izacard et al., 2021), M3E-Large [^9] https://github.com/wangyuxinwhy/uniem, mE5 (multilingual-e5-large) (Wang et al., 2024c); and OpenAI’s retriever, text-embedding-ada-002101010https://platform.openai.com/docs/guides/embeddings.
5.2. Main Results
@@
(1) BM25 remains highly competitive in specific retrieval scenarios. As a lexical retriever, it ranks documents based on TF-IDF matching scores calculated between queries and documents. Despite underperforming on the overall CMIRB benchmark, it displays strong results in tasks like medical news retrieval (78.9 vs. 73.33 for BGE) and medical post retrieval (66.95 vs. 67.13 for BGE). This can be attributed to the higher keyword overlap in these datasets, which aligns with BM25’s strengths.
(2) No single retrieval model achieves optimal performance across all tasks. PEG and GTE each deliver the best performance on four datasets, while BGE and mE5 each excel in achieving the top results on one dataset. Dense models with better performance often utilize contrastive learning, pretraining on large-scale unlabeled data followed by fine-tuning on labeled data. Variations in training data distribution influence model effectiveness across different datasets, suggesting the need for specialized approaches.
(3) SL-HyDE consistently outperformed HyDE across all ten datasets. While HyDE shows slight overall improvements over BGE, it excels in medical knowledge retrieval but underperforms in medical consultation tasks. This discrepancy could be due to LLM’s stronger handling of encyclopedia-type knowledge compared to the nuanced domain of patient-doctor consultations. In contrast, SL-HyDE achieved improvements over HyDE in all tasks, owing to its self-learning mechanism, which effectively enhances medical knowledge integration within both the generator and the retriever.
5.3. Performance Analysis
The core components of SL-HyDE are the generative LLM and the retrieval model. In this section, we explore the effects of varying these components.
Effect of Different LLMs   In Table 3, we present SL-HyDE’s performance with alternative LLMs fine-tuned, such as ChatGLM3-6B (Team et al., 2024) and Llama2-7b-Chat (Touvron et al., 2023), while keeping BGE as the retriever.
Both models demonstrate performance improvements under the SL-HyDE framework compared to HyDE. For instance, we observe a 4.65% improvement with ChatGLM3 and an 8.23% improvement with the Llama2 model. However, for Llama2, HyDE shows a slight decline compared to BGE. This is likely due to the fact that the pseudo-documents generated by the English-based Llama2 contained English content, which the downstream BGE retriever struggled to encode effectively. After fine-tuning, SL-HyDE improves by approximately 8%, attributed to both the reduction of English content and the enhanced retriever’s ability to encode medical knowledge, illustrating SL-HyDE’s adaptability across different LLMs.
Table 3. Performance of different generators.

### v2

@@ 3.3 SL-HyDE Training
Notably, we intentionally avoid using the true target document as the output label because the generator’s primary role is to craft a hypothetical document that aids the retriever in locating it. Expecting the generator to replicate the exact target document itself would be overly demanding and unrealistic.
@@
Given that not all hypothetical documents generated by the generator are equally effective for retrieval, we leverage the retriever $\mathcal{M}_{r}$ to select the most optimal one. Specifically, the generator $\mathcal{M}_{g}$ creates $L$ hypothetical documents for a given query. Each hypothetical document $d^{\prime}_{i}$ is used to retrieve documents from the corpus, and we record the rank $r_{i}$ of the true target document $d$. The pseudo-document with the highest retrieval quality (the lowest $r_{i}$) is selected:
$r_{i}=\mathrm{rank}(d,\mathrm{sort}(s(d^{\prime}_{i},D)),i=1,...,L,$ (4)
$i^{*}=\mathrm{arg\ min^{L}_{i=1}}\ r_{i},\ d^{*}=d^{\prime}_{i^{*}}.$ (5)
This process yields a collection of question-answer pairs in the form of $(q,d^{*})$, functions as the question and the generated document as the corresponding answer. The generator is subsequently trained via supervised fine-tuning on the resulting dataset $D_{llm}=\{(q,d^{*})|q\in Q\}$. The standard supervised fine-tuning (SFT) loss is computed as:
$\mathcal{L}_{\mathrm{slg}}=-\sum\nolimits_{q\in Q}\sum\nolimits_{t}\mathrm{log}\ \mathcal{M}_{g}(d^{\prime}_{t}|d^{\prime}_{\textless t},q).$ (6)
Interestingly, the self-learning generator is trained without relying on supervision signals from labeled medical data. Instead, it is based on unlabeled corpora and employs the generator’s text generation alongside the retriever’s ranking function to construct high-quality question-answer pairs tailored for hypothetical document generation.
Self-Learning Retriever. Given a passage $d$ from the corpus $D$ and its corresponding query $q$, the pair $(q,d)$ naturally forms the labeled query-document data required for retriever fine-tuning. However, since SL-HyDE retrieves the target document by encoding both the query and a generated hypothetical document when inference, we explore a triplet $(q,d^{\prime};d)$ as the labeled data for retriever training. This approach effectively aligns the training data format with that of the inference stage, thereby enhancing consistency and bridging the gap between training and deployment.
@@
To achieve this, we utilize the fine-tuned generator $\mathcal{M}_{g}^{t}$ from the previous stage to generate hypothetical documents for all queries, constructing a labeled fine-tuning dataset $D_{emb}=\{(q,d^{\prime};d)|q\in Q\}$. Following previous research Li et al. (2023b); Xiao et al. (2024), we further increase the training data complexity through hard negative mining. Specifically, a retriever is used to identify difficult negative samples from the original corpus $D$ through an ANN-based sampling strategy Xiong et al. (2020), resulting in a hard negative dataset:
$D^{-}=\mathrm{ANN}(\mathcal{M}_{r}(q,d^{\prime}),\mathcal{M}_{r}(D)).$ (7)
@@
In addition to the negatives mined from the corpus, we also incorporate in-batch negatives. Contrastive learning loss is then applied for the supervised fine-tuning of the retriever, with the objective function formulated as follows:
$\mathcal{L}_{\mathrm{slr}}=\mathrm{min.}\sum_{(q,d)}-\mathrm{log}\ \frac{e^{s(q,d)/\tau}}{e^{s(q,d)/\tau}+\sum\nolimits_{B\cup D^{-}}e^{s(q,d^{-})/\tau}},$ (8)
@@
where $\tau$ is the temperature coefficient, and $B$ represents the negative samples in a batch. The score $s(q,d)$ incorporates the generated document, as described in Equation 1.
At this stage, we can obtain a retriever equipped with medical domain knowledge that is coherently adapted to the characteristics of retrieval queries, incorporating hypothetical documents. In SL-HyDE, the generator and retriever are trained separately in a sequential manner, allowing each component to be optimized with the most appropriate supervision signal available at its respective training phase.
3.4 SL-HyDE Inference
As illustrated in Figure 1, the inference stage of SL-HyDE introduces a hypothesis generation step prior to conventional retrieval. Specifically, the input query $q$ is first rewritten by a fine-tuned generator $\mathcal{M}_{g}^{t}$ to produce a pseudo-document $d^{\prime}$, as defined by the following equation:
$d^{\prime}=\mathcal{M}_{g}^{t}(q,\mathrm{prompt}).$ (9)
The prompt is a manually designed instruction tailored to the requirements of each task. Detailed formulations of the prompts used in our experiments are provided in Appendix A.2.
To better fuse the documents, we sample $N$ documents from the hypothetical documents. Subsequently, a tuned retriever $\mathcal{M}_{r}^{t}$ is used to encode these documents into an embedding vector $v_{q}$:
$v_{q}=\frac{1}{N+1}[\sum_{k=1}^{N}\mathcal{M}_{r}^{t}(d^{\prime}_{k})+\mathcal{M}_{r}^{t}(q)].$ (10)
Then, the inner product is computed between $v_{q}$ and all document vectors:
$s(q,d)=<v_{q},\mathcal{M}_{r}^{t}(d)>,\forall d\in D.$ (11)
This vector identifies a neighborhood in the corpus embedding space, from which similar real documents are retrieved based on vector similarity.
4 CMIRB Benchmark
4.1 Overview
Figure 2: An overview of CMIRB.
The CMIRB benchmark is a specialized multi-task dataset designed specifically for medical information retrieval. As shown in Figure 2, it comprises five different tasks. Medical knowledge retrieval task: Retrieve relevant medical knowledge snippets from textbooks or encyclopedias based on a given medical query. Medical consultation retrieval task: Extract relevant doctor’s responses to online medical consultation questions posed by patients. Medical news retrieval task: Focus on retrieving news articles that address queries related to COVID-19. Medical post retrieval task: Retrieve the content of a forum post corresponding to its title. Medical literature retrieval task: Retrieve abstracts of cited references based on a medical title or find a similar paper based on the given medical paper.
4.2 Data Construction
The CMIRB benchmark integrates 10 datasets, including several existing resources: MedicalRetrieval Long et al. (2022), CmedqaRetrieval Qiu et al. (2022), and CovidRetrieval Qiu et al. (2022), covering patient-doctor consultations and COVID-19-related news retrieval.
In addition, we construct several datasets by combining existing query resources with curated medical corpora. MedExam pairs questions with textbook passages from MedQA Jin et al. (2021). DuBaike uses queries from DuReaderHe et al. (2017) and documents collected from Baidu Baike pages [^1] https://baike.baidu.com/. We also curate two datasets from the medical website DingXiangYuan [^2] https://dxy.com/. DXYDisease focuses on structured disease-related Q&A, while DXYConsult captures richer patient-doctor dialogues that include symptom descriptions, medication history, and diagnostic queries. We curate IIYiPost by crawling posts from the IIYi forum [^3] https://bbs.iiyi.com/.
@@ 4.2 Data Construction
Finally, CSLCite and CSLRel are constructed based on the CSL dataset Li et al. (2022), targeting different literature retrieval scenarios. CSLCite uses journal titles as queries and their cited references from WanFangMedical [^4] https://med.wanfangdata.com.cn/ as documents, while CSLRel pairs each paper with the most relevant similar paper recommended by the platform.
To ensure quality, we apply ChatGPT to exclude non-medical content and low-quality query-document pairs. Additional query-document matching is performed for MedExam and DuBaike to ensure content relevance. Full details are provided in the Appendix B.1. Table 1 summarizes dataset statistics, revealing broad variability in query and document length, ranging from short titles to long passages, ensuring the benchmark’s diversity and practical relevance.
@@
Table 1: Statistics of datasets in CMIRB.
@@
Table 2: Performance of various Retrieval models on nDCG@10. The first part shows ten base retrieval models, and the second shows retrieval models enhanced by hypothetical documents. denotes the result outperforms baseline models in t-test at $p<0.05$ level.
5 Experiments
5.1 Experimental Setup
Implementation Details. We sample 10,000 documents from the Huatuo26M_encyclopedia dataset as the unlabeled corpus. In our training framework, we utilize Qwen2-7B-Instruct Yang et al. (2024) as the generator and BGE-Large-zh-v1.5 Xiao et al. (2024) as the retriever. Unless otherwise stated, all experiments are conducted under this Qwen+BGE configuration. Model training and evaluation are conducted on up to 5 NVIDIA A100 GPUs, each equipped with 40GB of memory. For fine-tuning the LLM, we employ the AdamW optimizer Loshchilov (2017) in conjunction with a cosine learning rate scheduler. Training is executed for 1 epoch with a learning rate of 1e-5 and a batch size of 2. We set 200 warmup steps and configure the LoRA rank to 8. Retriever fine-tuning also uses the AdamW optimizer, with a linear decay schedule and an initial learning rate of 1e-5. The batch size per GPU is set at 4, and the maximum input sequence length is limited to 512. We apply a temperature of 0.02 and mine 7 hard negatives for each query to enhance training difficulty.
Evaluation Settings. For simplicity, we employ the LLM to generate a single hypothetical document for each query. The retrieval model embeds all queries, hypothetical documents, and corpus documents, with similarity scores calculated using cosine similarity. Documents in the corpus are ranked for each query based on these scores, and nDCG@10 is used as the primary evaluation metric to assess retrieval effectiveness. We set the temperature of LLM to 0.7 and repeat five times with different random seeds.
Baseline Models. To comprehensively evaluate CMIRB, we select several popular retrieval models. These include lexical retriever BM25 Robertson et al. (2009); dense retrieval models such as Text2Vec-Large-Chinese Xu (2023), PEG Wu et al. (2023), BGE-Large-zh-v1.5 Xiao et al. (2024), GTE-Large-zh Li et al. (2023b), and Piccolo-Large-zh SenseTime (2023); multilingual retrievers like mContriever (masmarco) Izacard et al. (2021), M3E-Large Wang et al. (2023b), mE5 (multilingual-e5-large) Wang et al. (2024a); and text-embedding-ada-002 OpenAI .
5.2 Main Results
@@
(1) BM25 remains highly competitive in specific medical tasks. As a lexical retriever, it ranks documents based on TF-IDF matching scores calculated between queries and documents. Despite underperforming on the overall CMIRB benchmark, it displays strong results in tasks like medical news retrieval (78.9 vs. 73.33 for BGE) and medical post retrieval (66.95 vs. 67.13 for BGE). This can be attributed to the higher keyword overlap in datasets.
(2) No single retrieval model achieves optimal performance across all ten tasks. PEG and GTE each deliver the best performance on four datasets, while BGE and mE5 each excel in achieving the top results on one dataset. Dense models with better performance often utilize contrastive learning, pretraining on large-scale unlabeled data followed by fine-tuning on labeled data. Variations in training data distribution influence model effectiveness across different datasets, suggesting the need for specialized approaches.
(3) SL-HyDE consistently outperformed HyDE across all ten datasets. While HyDE shows slight overall improvements over BGE, it excels in medical knowledge retrieval but underperforms in medical consultation tasks. This discrepancy could be due to LLM’s stronger handling of encyclopedia-type knowledge compared to the nuanced domain of patient-doctor consultations. In contrast, SL-HyDE achieved improvements over HyDE in all tasks, owing to its self-learning mechanism, which effectively enhances medical knowledge integration within both the generator and the retriever, while also aligning the outputs of the two models.
5.3 Performance Analysis
Effect of Different Generators. In Table 3, we present SL-HyDE’s performance with alternative fine-tuned LLMs as the generator, such as ChatGLM3-6B Team et al. (2024) and Llama2-7b-Chat Touvron et al. (2023).
Both models demonstrate performance improvements under SL-HyDE compared to HyDE. For instance, we observe a 4.65% improvement with ChatGLM3 and an 8.23% improvement with the Llama2 model. However, for Llama2, HyDE shows a slight decline compared to BGE. This is likely due to the fact that the pseudo-documents generated by the English-based Llama2 contained English content, which the downstream BGE retriever struggled to encode effectively. After fine-tuning, SL-HyDE improves by approximately 8%, attributed to both the reduction of English content and the enhanced retriever’s ability to encode medical knowledge, illustrating SL-HyDE’s adaptability.